OpenStudy (anonymous):

If (x + 2) is a factor of x3 − 6x2 + kx + 10, k =

OpenStudy (anonymous):

help @prowrestler

OpenStudy (anonymous):

@Spyro517 @slayer8001

OpenStudy (anonymous):

Let f(x) = x^3 - 6x^2 + kx + 10 If (x + 2) is a factor then f(-2) = 0 (-2)^3 - 6(-2)^2 + k(-2) + 10 = 0 -8 - 24 -2k + 10 = 0 -2k - 22 = 0 -2k = 22 k = -11